### v1

# Electricity Price Forecasting: Bridging Linear Models, Neural Networks and Online Learning

Abstract
Precise day-ahead forecasts for electricity prices are crucial to ensure efficient portfolio management, support strategic decision-making for power plant operations, enable efficient battery storage optimization, and facilitate demand response planning. However, developing an accurate prediction model is highly challenging in an uncertain and volatile market environment. For instance, although linear models generally exhibit competitive performance in predicting electricity prices with minimal computational requirements, they fail to capture relevant nonlinear relationships. Nonlinear models, on the other hand, can improve forecasting accuracy with a surge in computational costs. We propose a novel multivariate neural network approach that combines linear and nonlinear feed-forward neural structures. Unlike previous hybrid models, our approach integrates online learning and forecast combination for efficient training and accuracy improvement. It also incorporates all relevant characteristics, particularly the fundamental relationships arising from wind and solar generation, electricity demand patterns, related energy fuel and carbon markets, in addition to autoregressive dynamics and calendar effects. Compared to the current state-of-the-art benchmark models, the proposed forecasting method significantly reduces computational cost while delivering superior forecasting accuracy (12-13% RMSE and 15-18% MAE reductions). Our results are derived from a six-year forecasting study conducted on major European electricity markets.
Abstract
1 Introduction
The day-ahead electricity market operates through daily auctions submitted before noon, where prices are determined for electricity to be delivered on the following day ([neuhoff2016intraday]). It is the most important market in electricity trading and the main instrument for decision making ([m]). Accurate price forecasts support power plant operators in making optimal scheduling decisions for the following day. This is particularly important for units with technical constraints such as minimum downtime requirements and start-up costs. Precise forecasts are also essential for submitting competitive bids and, in turn, generating higher profits. They also play a key role in battery storage optimization by enabling operators to determine the most profitable charge–discharge strategy. Accurate forecasts are not only beneficial to supply-side decisions, but they are also crucial for demand-side management planning, such as designing cost-efficient electric vehicle charging strategies. In these decision-making settings, the time required to generate the forecast is equally important as its accuracy, since the market participants must use all available information while meeting strict deadlines. This underscores the importance of short-term predictive models with high precision within strict time constraints, which is the central objective of this paper.
The day-ahead forecast models can be classified into two main categories: Linear models and non-linear ones. Linear models with an autoregressive structure, where the lags of the price are used as predictors, were popular ([contreras2003arima]; [zhou2006electricity]). Later, the inclusion of exogenous variables such as solar and load along the lags has improved accuracy ([conejo2005forecasting]). More recently, shrinkage estimation methods such as least absolute shrinkage and selection operator (LASSO) ([tibshirani1996regression]) and Elastic Net ([zou2005regularization]) have demonstrated strong predictive accuracy thanks to their ability to select relevant features ([lear]; [ziel2016forecasting], [ziel2018day]). Linear models are generally effective in predicting electricity prices and are highly valued for their interpretability and fast computation. However, due to their basic structure, they may still miss important non-linear relationships, such as the interaction between variables which can lead to less accurate forecasts.
Those interactions, on the other hand, could be captured by non-linear models, which generally consist of many classes, particularly from machine learning models such as decision tree learners, like random forests, and gradient boosting machines; and neural networks. However, we will focus here on the deep learning methods, given their rapidly growing popularity in electricity price forecasting ([a]). The multilayer perceptron (MLP) is one of the simple architectures that has already been employed to generate forecasts a few years back ([szkuta2002electricity]; [catalao2007short]). Later, a more sophisticated architecture, like a denoising autoencoder, originally popular in natural language processing, is employed for price forecasting ([wang2016short]). Additionally, Recurrent neural network (RNN) is also widely employed due to its structure to capture temporal dependencies in time series data ([dnn]; [ugurlu2018electricity]; [chinnathambi2018deep]). Note that the deep learning methods usually outperform the linear models in terms of accuracy but fails in terms of computational cost ([a]). To address this trade-off, we build a hybrid model that combines the strengths of both approaches while following the standard framework checklist proposed by [a].
We propose a simple neural network architecture that combines input and output through two components. The first component is a generalized linear model that connects the input to the output directly, while the second is an MLP that connects the input to the output through one hidden layer with a nonlinear activation function. This architecture was first introduced by [google] to improve application recommendations. Thanks to this structure, the model can both learn the explicit linear relationship and the hidden nonlinear interactions. In addition to this unique architecture, the model incorporates all key drivers of electricity price formation, including renewable generation, electricity demand, fuel and carbon prices, autoregressive dynamics, and calendar effects. This ensures that both fundamental and seasonal relationships are fully captured. It is worth emphasizing that the literature is rich in hybrid models that combine different algorithms to improve accuracy in electricity price forecasting ([kim2015short]; [ben2018forecasting]; [bento2018bat]; [bisoi2020short]; [gomez2023electricity]; [cu2025time] ). However, none of the existing studies have employed the proposed hybrid model.
To further improve accuracy, an ensemble method is implemented to average the individual forecast models. Combining the predictions of different models has been shown to enhance precision as compared to individual forecasts in many studies ([raviv2015forecasting], [nowotarski2016combine], [maciejowska2020pca]). The arithmetic mean is the simplest averaging technique, such that equal weight is assigned to each individual forecast ([stock2004combination]). In this work, we select fully adaptive Bernstein Online Aggregation (BOA) proposed by [wintenberger2017optimal] as an ensemble method. This approach assigns dynamic weights to individual models based on their recent performance. While fully adaptive BOA was previously used to predict the electricity load ([de2023adaptive]; [hirsch2024online]; [himych2024adaptive]), to the best of our knowledge, it has not yet been used for electricity price forecasting. .
In addition to improving accuracy, keeping the computational time at an efficient level is a key objective of this work. To this end, we implement a novel partial online learning approach designed to minimize the computational burden. This method employs information from the previous iteration to reduce the training time of the model. While online learning algorithms were formerly employed in the electricity price literature, for instance [hirsch2025online] used it to generate a probabilistic price forecast, there is no paper in the literature that has applied the non-parametric online learning that we introduced to forecast electricity prices.
Finally, the performance of the proposed models is compared to two well-established and performing models in electricity price forecasting ([a]), which are Lasso Estimated AutoRegressive (LEAR) ([lear]) and the Deep Neural Network (DNN) ([dnn]). We implement a forecasting study using the suggested methodology for two major European day-ahead electricity markets: the German–Luxembourg and the Spanish markets, with the former being the largest in Europe. The results demonstrate that our approach yields a powerful model that significantly outperforms the benchmark models in terms of accuracy and computational time.
The main contributions of this work can be summarized as follows:
  1. 1.
    We provide the first implementation of the proposed linear–nonlinear hybrid architecture for day-ahead electricity price forecasting.
  2. 2.
    We are the first to apply a fully adaptive BOA framework for model ensembling and forecast combination.
  3. 3.
    We introduce a new non-parametric partial online learning approach suitable for both linear and neural network models, enabling fast and efficient learning.
  4. 4.
    We conduct an extensive multi-market empirical study using more than six years of hourly data, demonstrating the robustness and transferability of the proposed methodology across different market structures.
  5. 5.
    The results shows that the proposed models consistently outperform the current state-of-the-art approaches in terms of both accuracy and computational efficiency across two major European markets, highlighting the general applicability of the framework.
Besides this introduction, the manuscript consists of five more sections. Section 2 provides an overview of the data. Section 3 presents a general framework of model design. The forecast study design and the evaluation measures are described in Section 4, while results are presented in Section 5. Section 6 summarizes the findings and provides an outlook for future work.
2 Data
The datasets used in this project are mainly extracted from the website https://www.entsoe.eu of the European Network of Transmission System Operators for Electricity (ENTSO-E), which is a platform that provides complete information on the European electricity market. We consider data from 2018-01-10 to 2025-01-14 for the two largest European electricity markets: German-Luxembourg [^1] Note that in October 2018 the Austrian-German-Luxembourg price zone got split. and Spain. Thus, each of the datasets contains 2298 days of hourly observations (55153 hours) of the main variables for the electricity market.
For each of the considered datasets, the objective variable is the hourly day-ahead electricity price, which we denote $P_{d,h}$ for the price at day $d$ and hour $h$ where $h=0,\ldots,23$. We also consider typical external data, particularly day-ahead forecasts from renewable energy sources, solar ($\text{Solar}_{d,h}$), wind onshore ($\text{WindOn}_{d,h}$) and wind offshore ($\text{WindOff}_{d,h}$), as well as day-ahead load forecasts ($\text{Load}_{d,h}$). Note that wind offshore is only available on the German-Luxembourg dataset but not in the Spanish one, as Spain currently has no large operational offshore wind generation facilities ([cerda2025present]). In addition to these fundamental variables, both datasets incorporate four electricity market-related commodity price time series. These are closing prices of short-term commodity futures, of TTF natural gas price ($\text{NGas}_{d}$), Brent oil price ($\text{Oil}_{d}$), Rotterdam Coal price ($\text{Coal}_{d}$) and carbon emission prices by the European Emission Allowances ($\text{EUA}_{d}$). The commodity prices are collected from refinitiv datastream ([refinitiv_datastream]).
The fuel-related prices are the same for both datasets, since they are not country-specific. The missing values in these commodity variables are the only missing values in the German-Luxembourg dataset. Those missing are replaced with the last available values, which are in line with energy market literature and correspond to the martingale assumption on these price time series. In contrast, the Spanish dataset contains additional missing values in the renewable and load day-ahead variables. Those missing are imputed by predictions obtained from simple regressions of the day-ahead variable on its corresponding actual variables. Lastly, given the distortions caused by daylight saving time, the missing hour on the last Sunday of March is imputed by the means of the adjacent hours, and the duplicated values on the last Sunday in October are replaced by their average.
The time series in Figure 1 provides a comprehensive overview of the changing dynamics of the electricity price in the German-Luxembourg bidding zone. Note that, in the main text, we focus on the data and result visualizations for this market, as it is the largest electricity market in Europe. The plots for the Spanish market are provided in the appendix.
Figure 1: Time Series of Day-Ahead Electricity Prices in German-Luxembourg
Examining this price series in Figure 1, one can see a high fluctuation, with a notable peak around August 2022, likely driven by the energy crisis following the Russia-Ukraine conflict. Subsequently, prices gradually return to their typical interval, with some peaks and troughs that can be attributed to variations in demand and renewable generation. Another distinctive feature highlighted in Figure 1, which characterizes the German-Luxembourg market and is not common across all electricity markets, is the existence of negative prices. These can arise from subsidy schemes (renewable generators keep producing regardless of the market price) or production constraints (costs of keeping the power plant on could be smaller than shutting down and restarting it later).
Figure 2 depicts the fundamentals and commodity variables employed as regressors to generate the forecast for the German-Luxembourg market. The bottom subplot highlights the surge in natural gas prices during the energy crisis. At the same time, the sharp increase in coal prices is attributed to the extensive use of coal-fired power during this crisis to compensate for the shortage of natural gas ([huang2023research]). These developments likely contributed to the rise of electricity prices shown in Figure 1.
Figure 2: Day-Ahead Forecasts of Load and Renewable Generation, and Commodity Prices in the German-Luxembourg Electricity Market
Fundamental variables in Figure 2 show a clear annual seasonal pattern. Solar generation is low in winter and relatively high during the rest of the year. In contrast, electricity load is high during winter with a noticeable trough during the holiday season. In addition to yearly seasonality, electricity load exhibits weekly seasonality, which is not very apparent in the above Figure. The load is low during off-peak hours and weekends, which is usually associated with a decrease in electricity prices.
3 Model Design
Now we proposes and compares various models designed explicitly for day-ahead electricity price forecasting models.We examine two main configurations, linear and combined structure. Afterwards we explain the proposed ensembling method in detail.
3.1 Linear Networks
This architecture consists of an input layer and an output layer with 24 neurons, where each neuron represents the price of electricity at a specific hour. This setup is employed as the basis for two variants, referred to as the Reduced Linear and Full Linear models. The Reduced Linear predicts the hourly electricity prices using regressors of the specific predicted hour. For instance, to predict hour zero (midnight to one o’clock in the morning), the model uses only the lagged prices and fundamentals corresponding to hour 0, along with time-invariant variables such as weekday dummies and commodity prices. The Reduced Linear model is expressed as follows:
$P_{d,h}$ $=\beta_{h,0}+\beta_{h,1}P_{d-1,h}+\underbrace{\beta_{h,2_{2}}P_{d-2,h}+\beta_{h,2_{3}}P_{d-7,h}}_{\text{$=\bm{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,h}$}}+\beta_{h,3}P_{d-1,23}+$
$\quad\underbrace{\beta_{h,4_{1}}\text{Solar}_{d,h}+\beta_{h,4_{2}}\text{WindOn}_{d,h}+\beta_{h,4_{3}}\text{WindOff}_{d,h}+\beta_{h,4_{4}}\text{Load}_{d,h}}_{\text{$=\bm{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,h}$}}+$
$\quad\underbrace{\beta_{h,5_{1}}\text{Oil}_{5-2}+\beta_{h,5_{2}}\text{Coal}_{d-2}+\beta_{h,5_{3}}\text{EUA}_{d-2}+\beta_{h,5_{4}}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\quad\underbrace{\beta_{h,6_{1}}\text{Mon}_{d}+\beta_{h,6_{2}}\text{Sat}_{d}+\beta_{h,6_{3}}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\epsilon_{d,h}$ (1)
  • $\mathbf{P}_{\cdot,h}=\big(P_{d-2,h},\;P_{d-7,h}\big)^{\prime}\quad\text{(lag 2 and 7 of price)},$ which includes price information at the same hour, $h$, 2 and 7 days ago.
  • $\mathbf{Fun}_{d,h}=\big(\text{Solar}_{d,h},\;\text{WindOn}_{d,h},\;\text{WindOff}_{d,h},\;\text{Load}_{d,h}\big)^{\prime}\quad\text{(Fundamentals)},$ includes fundamental day-ahead forecast information at hour $h$ from load, wind and solar power
  • $\mathbf{Com}_{d-2}=\big(\text{Oil}_{d-2},\;\text{Coal}_{d-2},\;\text{EUA}_{d-2},\;\text{NGas}_{d-2}\big)^{\prime}\quad\text{(Commodities)},$ represents a vector of the most recent commodity prices from natural gas, coal, oil and carbon emissions.
  • $\mathbf{Cal}_{d}=\big(\text{Monday}_{d},\;\text{Satuday}_{d},\;\text{Sunday}_{d}\big)^{\prime}\quad\text{(Calendar dummies)},$
and the error term $\epsilon_{d,h}$. The coefficients $\beta_{d,j}$’s represent the model parameters to be estimated or trained. Commodity prices are modeled using information available on day d2d-2, since at the time of bidding (before noon on day d1d-1) the stock closing prices of commodities are not yet available.
In addition to the Reduced Linear model, we consider a Full Linear model, which uses regressors from all hours rather than restricting them to the target predicted hour $h$. For example, when predicting the price for hour $h=0$, the model incorporates lagged prices and fundamentals from hours 0 through 23 as predictors, together with regressors that do not vary within the day. The following equation provides the Full Linear model specification:
$P_{d,h}$ $=\beta_{h,0}+\sum_{l=0}^{23}\beta_{h,l,1}P_{d-1,l}+\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,2_{2}}P_{d-2,l}+\beta_{h,l,2_{3}}P_{d-7,l}\right)}_{\text{$=\bm{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,l}$}}+$
$\qquad\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,4_{1}}\text{Solar}_{d,l}+\beta_{h,l,4_{2}}\text{WindOn}_{d,l}+\beta_{h,l,4_{3}}\text{WindOff}_{d,l}+\beta_{h,l,4_{4}}\text{Load}_{d,l}\right)}_{\text{$=\bm{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,l}$}}+$
$\qquad\underbrace{\beta_{h,5_{1}}\text{Oil}_{d-2}+\beta_{h,5_{2}}\text{Coal}_{d-2}+\beta_{h,5_{3}}\text{EUA}_{d-2}+\beta_{h,5_{4}}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\qquad\underbrace{\beta_{h,6_{1}}\text{Mon}_{d}+\beta_{h,6_{2}}\text{Sat}_{d}+\beta_{h,6_{3}}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\nu_{d,h},$ (2)
where $\nu_{d,h}$ is the error terms. The regressor $P_{d-1,23}$ is omitted from equation (2), as it is already included in the term $\sum_{l=0}^{23}\beta_{h,l,1}P_{d-1,l}$.
3.2 Non-Linear Networks and Combined structures
The second main architecture in this project is a more elaborate network that, in addition to input to the output layer, employs a hidden layer equipped with the Leaky Relu activation function to capture potential nonlinearities. Two specifications of this architecture are considered: the MLP with Reduced Linear and the MLP with Full Linear. The MLP with Reduced Linear is the sum of the linear predictor of equation (1) and a nonlinear part $\mathrm{NN}(\theta)$:
$P_{d,h}$ $=\;\beta_{h,0}+\beta_{h,1}P_{d-1,h}+\mathbf{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,h}+\beta_{h,3}P_{d-1,23}+\mathbf{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,h}+$
$\quad\mathbf{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}+\mathbf{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}+\mathrm{NN}(\mathbf{\theta})+\xi_{d,h},$ (3)
$\mathrm{NN}(\mathbf{\theta})(\mathbf{x})\;=\;\mathbf{W}^{(2)}\,\sigma_{\alpha}\bigl(\mathbf{W}^{(1)}\,\mathbf{x}\;+\;\mathbf{b}^{(1)}\bigr)\;+\;\mathbf{b}^{(2)},\;\;\;\;\sigma_{\alpha}(z)\;=\;\begin{cases}z,&z\geq 0\\ \alpha\,z,&z<0\end{cases},$
where $\mathbf{x}\in\mathbb{R}^{m}$ is the input vector with $m$ regressors, $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\mathbf{b}^{(1)},\mathbf{b}^{(2)}$ are the trainable weight matrices and bias vectors, respectively. $\mathbf{\theta}$ denotes the collection of all trainable parameters, and $\alpha=0.01$, is the leakage factor in the Leaky ReLU activation function.
Figure 3 illustrates the architecture of MLP with Reduced Linear Model. The green neurons represent the input layer. The hidden layer is defined by the blue $n$ neurons, where the number of neurons $n$ is a hyperparameter to be tuned. The output layer consists of 24 neurons, where each neuron provides the forecast for a specific hour of the day, is highlighted in orange. The blue edges linking the input and hidden layers correspond to the weight matrix $\mathbf{W}^{(1)}$, while the remaining blue connections are saved in $\mathbf{W}^{(2)}$. For simplicity, the bias vectors $\mathbf{b}^{(1)}$ and $\mathbf{b}^{(2)}$ are not shown in Figure 3. Note that the rectangular neurons represent vectors rather than scalars. Thus, any connection from a rectangular node to another neuron does not represent a single weight, but a vector of weights. For instance, the blue edges connecting $\mathbf{Com}_{d-2}$ to the hidden neuron $\mathbf{H}_{1}$ illustrate four learnable weights, since $\mathbf{Com}_{d-2}$ is a 4-dimensional vector.
Figure 3: MLP with Reduced Linear Model architecture to forecast the hourly electricity price
The red edges directly linking the input to the output, referred to as skip connection, represents the model’s linear predictor. The red solid lines connect to the variables that vary for each hour, whereas the dashed red lines correspond to variables that are constant throughout the day, such as fundamentals. However, an exception arises for the covariate $P_{d-1,23}$, which is constant across hours and also serves as the lagged price for predicting $P_{d,23}$. Accordingly, the skip connection for hours $P_{d,0}$ through $P_{d,22}$ includes 15 regressors, while the skip connection for $P_{d,23}$ contains only 14 covariates. It is worth noting that the linear relationships are not only learned through skip connection since the MLP component of the model can learn linear dependencies as well, which results in identifiability issues. However, one can include the orthogonalization cell proposed by [rugamer] to restore the interpretability of the linear components.
MLP with Full Linear model shares the same hidden layer design as the MLP with Reduced Linear model, which is defined as $\mathrm{NN}(\mathbf{\theta})$ in equation (3). Yet differs in how the input-to-output connection is structured, as this linear component corresponds to that of the Full Linear model. The model equation is expressed in the following form:
$P_{d,h}$ $=\;\beta_{h,0}+\sum_{l=0}^{23}\beta_{h,l,1}P_{d-1,l}+\sum_{l=0}^{23}\mathbf{\beta}_{h,l,2}^{\prime}\mathbf{P}_{\cdot,l}+\sum_{l=0}^{23}\mathbf{\beta}_{h,l,4}^{\prime}\mathbf{Fun}_{d,l}+$
$\quad\mathbf{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}+\mathbf{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}+\mathrm{NN}(\mathbf{\theta})+\mu_{d,h}.$ (4)
Hence, the MLP with Full Linear model employs 175 regressors to forecast each hour.
In addition to the four models introduced above, we develop three further variants. The first is the plain MLP model, which consists of an input layer, a hidden layer, and an output layer, without any skip connections such that the coefficients of the linear component in equation (3) are set to zero ($\beta_{h,0}=\beta_{h,1}=\beta_{h,3}=\mathbf{\beta}_{h,4}^{\prime}=\mathbf{\beta}_{h,5}^{\prime}=\mathbf{\beta}_{h,6}^{\prime}=0$). The second and third are the Reduced Linear with OLS and the MLP with Reduced Linear and OLS models. These are structurally similar to the Reduced Linear and the MLP with Reduced Linear models, respectively, but differ in their weight initialization: the OLS-based variants use the coefficients of a simple linear model estimated via OLS as initial weights, whereas the non-OLS counterparts initialize weights randomly.
It is important to highlight that each of the above models is implemented identically to predict electricity prices in the German-Luxembourg and Spain bidding zones. The only exception is that $\beta_{h,4_{3}}$ is set to zero in the Spanish case due to the absence of offshore wind generation data. Furthermore, all the regressors are standardized to enhance training stability.
3.3 Ensembling
Ensemble learning is a method where the forecasts of different models, referred to as experts, are combined with the goal of improving accuracy. Each of the models is multiplied by a specific weight and then aggregated to generate a single forecast such that the weights sum up to one:
$\hat{P}_{d,h}^{\text{ens}}=\sum_{k=1}^{K}w_{k,d,h}\hat{P}_{k,d,h}$
where $\hat{P}_{d,h}^{\text{ens}}$is the ensemble forecast at day $d$ and hour $h$ , $\hat{P}_{k,d,h}$ is the forecast of model $k$, $w_{k,d,h}$ is the corresponding weight, and $K$ is the number of models.
The simplest ensemble approach is to assign an equal weight to all models such that $w_{k,d,h}=\frac{1}{K}$. This weighting aggregation is employed, for example, by Random Forest ([breiman2001random]), where the forecasts of different trees (experts) are averaged. However, it is important to highlight that the strength of Random Forest lies not only in ensembling, but also in the construction of decorrelated trees and random feature selection [hastie2009elements, Ch. 15].
A more advanced ensemble method is the fully adaptive Bernstein Online Aggregation (BOA) method proposed by [wintenberger2017optimal]. The weight of this approach varies among experts and between time steps for an individual expert. It is expressed as follows
$w_{k,d,h}=\frac{\eta_{k,d-1,h}\,\exp\!\big(-\eta_{k,d-1,h}R_{k,d-1,h}\big)}{\sum_{j=1}^{K}\eta_{j,d-1,h}\,\exp\!\big(-\eta_{j,d-1,h}R_{j,d-1,h}\big)},$
where $\eta_{k,d-1,h}$ is the adaptive learning rate for expert $k$, and $R_{k,d-1,h}$ is the cumulative corrected regret of expert $k$, computed up to day d1d-1 for hour $h$. The inclusion of $R_{k,d-1,h}$ inside the exponential ensures that experts who performed poorly in the past (up to day d1d-1) receive a smaller weight in day $d$, while larger weights are assigned to the best-performing experts in the past. Moreover, unlike classical ensemble methods such as Exponentially Weighted Average (EWA) ([vovk1990aggregating]), fully adaptive BOA employed the second-order refinement that stabilises the learning process.
4 Training and Evaluation Setup
4.1 Learning Algorithm
In a normal rolling window forecast study, the forecast is generated following this mechanism: data is divided into the test data with $P$ observations and training of size $D$, where D=TPD=T-P and $T$ is the sample size. The model is fitted on observations $t=1$ to $t=D$, and the forecast is generated on observation $t=D+1$, which is the first observation in the test data. In the next step, the model is fitted on observations $t=2$ to $t=D+1$, and the forecast is generated on the second observation of the test data, $t=D+2$, and so on.
The issue with this mechanism is that in each iteration, the model is refitted on the last $D$ observations with only a single additional observation compared to the previous iteration, which makes the procedure time-consuming. To reduce this computational burden, online learning is introduced such that instead of refitting all $D$ observations, the weights are updated in each iteration using the formerly estimated coefficients along the new observation. Online learning is applied in electricity price forecasting for both point and probabilistic predictions. For instance, [hirsch2024online] implements an algorithm that embeds online coordinate descent (OCD) or recursive least squares (RLS) into the Generalized Additive Models for Location, Scale and Shape (GAMLSS) framework to estimate regularized linear distributional models. Additionally, [yan2012online] employs an online/incremental Support Vector Regression (SVR) approach, which updates the model parameters as new data arrive, instead of retraining from scratch, to generate point forecasts of electricity prices.
This project, on the other hand, implements a partial online learning using a novel non-parametric approach depicted in Figure 4. In the first iteration, the model is initialized randomly and fitted using 10 observations. The resulting weights are then employed to forecast observation 11. The optimal weights from the first iteration are used as initialization of the model in the second iteration, and only the most recent five observations are used to refit the model. Similarly, in iteration 3, the optimal weights from iteration 2 are carried forward, and the model is retrained on the latest five observations. This process continues till the final observation $T$ is forecasted. This approach is constructed on the assumption that the model requires a larger sample in the first iteration to learn the data structure, while in subsequent iterations, fewer observations suffice, since the carried-over weights already contain information from past data. The initial 10 observations are referred to as the initial window size, whereas the subsequent five observations used in each iteration define the update window size. The name partial comes from the use of a couple of observations in the updated iterations, not the new observation only, which is the standard in the online learning algorithm.
Figure 4: Visualisation of the online learning algorithm
The window size visualization in Figure 4 serves only as a simple illustration of how the online learning procedure functions. However, there are other essential elements for the algorithm to work effectively, notably: the number of epochs, which determines the number of times the model sees the full dataset, and the learning rate, which controls the magnitude of parameter updates during training. The number of epochs in the initial iteration should generally be larger than the number of epochs in the updated windows. Since a higher number of epochs is associated with a larger number of weight updates, this ensures convergence toward a local minimum. However, for the subsequent iteration, the model starts in the neighborhood of the minimum thanks to initialization weights from the previous iteration, and then a small number of epochs is usually adequate to ensure convergence. A similar principle applies to the learning rate.
4.2 Study Design
The dataset is split into training, validation, and test sets, where the test data comprises the last two years of observations from 2023-01-16 to 2025-01-14, and the validation set includes the two prior years of observations. Training data, on the other hand, changes in every iteration. For instance, during the hyperparameter tuning phase, where the best hyperparameters are selected, the training set is the $D_{\text{init}}$ observations prior to 2021-01-16, which marks the first observation in the validation dataset. The second training set then comprises the $D_{\text{up}}$ observations preceding the second validation point, and so on, where $D_{\text{init}}$ is the window size of the initial window and $D_{\text{up}}$ is the window size of the updated window.
Both $D_{\text{init}}$ and $D_{\text{up}}$ are hyperparameters to be tuned, along with the number of neurons in the hidden layer and initial and updated learning rate. In addition, the penalty terms $\lambda_{1}$ and $\lambda_{2}$ in the loss function (as defined in the following equation) are also subject to tuning:
$L(\bm{\theta})=\ell(\mathbf{P})+\lambda_{1}\|\bm{\theta}\|_{2}^{2}+\lambda_{2}\|\bm{w}^{(\mathrm{out})}\|_{1}$
where $\ell(\mathbf{P})$ is L1 loss, $\|\bm{\theta}\|_{2}$ is the Euclidean norm of the weight vector $\bm{\theta}$, and $\|\bm{w}^{(\mathrm{out})}\|_{1}$ is norm of the output-layer weights. The use of this regularized loss function is necessary to avoid overfitting, which is particularly important given the large number of regressors. The last hyperparameter to consider for tuning is solely employed by the two OLS-based models. It is denoted as $\alpha$ and it determines the share of the OLS coefficient, $\bm{\hat{\beta}}_{\text{OLS}}$, to be used for initializing the skip connection of Reduced Linear with OLS and the MLP with Reduced Linear and OLS models. For example, if $\alpha=1$, the initialization is given directly by $\bm{\hat{\beta}}_{\text{OLS}}$, while if $\alpha=2$, then models will use $2\times\bm{\hat{\beta}}_{\text{OLS}}$ as initialisation.
The Tree-structured Parzen Estimator (TPE) optimization method, proposed by [TPE], is employed to select the combination of optimal values of the above hyperparameters. TPE is a Bayesian optimization algorithm that selects candidate hyperparameters by modeling the distribution of good and bad trials from past evaluations to minimize the objective function. Unlike random search, which samples random values within the proposed interval in each trial, TPE employs previous trials to guide the search toward more promising regions of the hyperparameter space. We implement this hyperparameter optimization algorithm using the Python package Optuna ([optuna]), where we set the number of trials to 500 and set the mean absolute error as an objective function.
On the other hand, there are other hyperparameters that we predetermine, such as the number of epochs, which we set to 60 for the initial iteration and 10 in the subsequent iterations, and the optimization algorithm, which we fix to Adaptive Moment Estimation (Adam) as it is simple and computationally efficient ([adam]).
During hyperparameter tuning, 500 different combinations of hyperparameters are created for each of the 7 model classes discussed in section 3. The combination that yields the smallest MAE is employed for fitting and generating the forecast in the test data. Additionally, for each model class $i$, where i=1,2,7i=1,2,...7 we select the ten models that yield the best ensemble performance using the following iterative procedure:
  • Start with the model $\mathcal{M}_{1}^{i}$ that achieves the lowest MAE among all 500 candidate models.
  • Form pairwise BOA ensembles of $\mathcal{M}_{1}^{i}$ with each of the remaining 499 models. Select $\mathcal{M}_{2}^{i}$, the model that together with $\mathcal{M}_{1}^{i}$ produces the lowest MAE.
  • Next, form BOA ensembles of the current set $\{\mathcal{M}_{1}^{i},\mathcal{M}_{2}^{i}\}$ with each of the remaining 498 models. Select $\mathcal{M}_{3}^{i}$, the model that yields the lowest MAE when added to the ensemble.
  • Continue this forward step selection process until an ensemble of 10 models has been constructed.
The resulting model for class $i$ will be denoted Name (BOA). For instance, the ensemble form resulting from the Reduced Linear architecture will be referred to as Reduced Linear (BOA). In addition to these seven class-specific ensembles, we also develop a general ensemble model, referred to as BOA all. This ensemble is built from the 500 models with the lowest MAE across all model classes. The same forward selection strategy is applied to identify the best combination of ten models. Note that all models, including individual and ensemble variants, are selected based on their performance on the validation dataset.
4.3 Benchmarks and Evaluation
In order to check the strengths of our models, two state-of-the-art models are selected as benchmarks. The first benchmark is Lasso Estimated AutoRegressive (LEAR), which is a high-dimensional linear benchmark for electricity price forecasting. LEAR coefficients are estimated using L1 regularization (LASSO) to enforce sparsity and prevent overfitting ([lear]). The second benchmark is the Deep Neural Network (DNN), which is a multilayer perceptron with input, output, and two hidden layers. DNN employs ADAM as an optimization algorithm for the weights and the Tree-structured Parzen Estimator algorithm to tune the hyperparameters ([dnn]).
Both LEAR and DNN are implemented using the Python library epftoolbox ([toolbox]) with a training window size of two years and the following regressors:
  • autoregressive lags of the price ($P_{d-1,h},P_{d-2,h},P_{d-3,h},P_{d-7,h}$),
  • day of the week dummies,
  • renewables as one regressor ($\text{RES}_{i,h}=\text{Solar}_{i,h}+\text{WindOn}_{i,h}+\text{WindOff}_{i,h}$) and $\text{load}_{i,h}$ for i=d,d1,d7i=d,d-1,d-7.
Finally, three accuracy metrics are selected to evaluate and compare the performance of all discussed models: the root mean square error (RMSE), the mean absolute error (MAE), and the relative mean absolute error (rMAE)
RMSE $=\sqrt{\frac{1}{24N_{d}}\sum_{d=1}^{N_{d}}\sum_{h=0}^{23}\big(\hat{P}_{d,h}^{\text{model}}-P_{d,h}\big)^{2}}$
MAE $=\frac{1}{24N_{d}}\sum_{d=1}^{N_{d}}\sum_{h=0}^{23}\left|\hat{P}_{d,h}^{\text{model}}-P_{d,h}\right|$
rMAE $=\frac{\frac{1}{24N_{d}}\sum_{d=1}^{N_{d}}\sum_{h=0}^{23}\left|\hat{p}_{d,h}^{\text{model}}-p_{d,h}\right|}{\frac{1}{24N_{d}}\sum_{d=1}^{N_{d}}\sum_{h=0}^{23}\left|\hat{p}_{d,h}^{\text{naive}}-p_{d,h}\right|}$
where $\hat{p}_{d,h}^{\text{model}}$ is the forecasted price of the specific model while ${p}_{d,h}$ is the real price. $N_{d}$ is the number of days in the test period which is set 730 days. $\hat{p}_{d,h}^{\text{naive}}$ is the forecast price of the following naive model:
[…]
For each of the above accuracy metrics, the model with the smallest value of the corresponding metric is considered the best-performing model. However, there is no guarantee that this apparent superiority is statistically significant and could simply result from random chance. Hence, along with the accuracy metrics, the Diebold-Mariano (DM) test is employed, where the DM test is a statistical test that compares the forecast accuracy of two forecasting models ([dm]). Let $\hat{P}_{d,h}^{(A)}$ and $\hat{P}_{d,h}^{(B)}$ be the generated forecast at day $d$ on hour $h$ using models $A$ and $B$, respectively. The forecast error of model $A$ is $\hat{e}_{d}^{(A)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(A)}-\sum_{h=0}^{23}{P}_{d,h}^{(A)}\right)$ and forecast error of model $B$ is $\hat{e}_{d}^{(B)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(B)}-\sum_{h=0}^{23}{P}_{d,h}^{(B)}\right)$.
The null hypothesis to compare the forecast accuracy of $A$ and $B$ is denotes as $E\left(\left|\hat{e}_{d}^{(A)}\right|-\left|\hat{e}_{d}^{(B)}\right|\right)=0,$ where $E(.)$ is the expectation. The null hypothesis implies that there is no difference in forecast accuracy between $A$ and $B$. The alternative hypothesis is expressed as: $E\left(\left|\hat{e}_{d}^{(A)}\right|-\left|\hat{e}_{d}^{(B)}\right|\right)<0,$ which suggests that the forecasts of model $A$ are statistically more accurate than those of $B$.
5 Results and Discussion
After selecting the optimal models based on their performance on each of the validation datasets, the forecast accuracy of those models is evaluated on the corresponding test data employing accuracy metrics and a statistical test. Figure 5 depicts the hourly RMSE of the main models, along with benchmark models for the German-Luxembourg market. Studying Figure 5, one can notice that all models follow the same pattern: perform well during the first and last hours of the day, whereas their performance deteriorates during the peak hours.
Figure 5: Hourly RMSE of selected models for German-Luxembourg bidding zone
During the first hour of the day, the LEAR has the lowest RMSE. However, its performance worsened afterwards; it even has the worst performance among all hours, starting from hour 15. The second benchmark model, which is DNN, performs poorly in most hours of the day. Except for the hour zero, MLP with Reduced Linear and MLP with Reduced Linear (BOA) generate the most accurate forecast compared to other models; they interchangeably have the lowest RMSE. For instance, MLP with Reduced Linear is the best performing model in hours 5 and 22, with values of 11.66 and 12.46, respectively, while MLP with Reduced Linear (BOA) has the lowest RMSE in hours 2 and the last hour.
It is apparent that the model performance changes from one hour to another. Hence, to select the best performing model, three overall accuracy measures are calculated and the results are presented in Table 1. The green cells in Table 1 correspond to the best performing model for the respective metric, while the red cells depict the worst performing ones.
Table 1: Comparative Forecast Accuracy in terms of RMSE (EUR/MWh), MAE (EUR/MWh) and rMAE of Competing Models
For both the German-Luxembourg and Spanish bidding zones, MLP with Reduced Linear (BOA) is the best performing model according to MAE and rMAE. On the other hand, when considering RMSE, the MLP with Reduced Linear model is the most accurate in the German–Luxembourg market, whereas the emphMLP with Full Linear model achieves the best performance in the Spanish market. It is apparent that all the developed models in this paper outperformed the state-of-the-art benchmark models in both markets.
The RMSE is reduced by 12-13% and the MAE by 15-18% when comparing the best models to the best-performing benchmark, DNN, in the respective market. Additionally, one can notice an overall decrease in both RMSE and MAE in the Spanish market compared to the same models in the German-Luxembourg market. This might be the result of the share of renewables in the electricity generation. [rintamaki2017does] confirmed that wind-powered generation increases the price volatility in Germany. On the other hand, [pereira2017effect] concluded that while wind increases volatility in Spain, the integration of hydro power plants decreases it. Finally, it is worth emphasizing that for the three measures, we can see a significant improvement from the MLP model to the MLP with Reduced Linear. For instance, the RMSE of MLP with Reduced Linear in the Spanish bidding zone is smaller than that of the MLP by 10%. This highlights the importance of the novel architecture that we proposed in improving accuracy compared to a neural network without a skip connection.
Figure 6 presents the heatmap of the p-value of the Diebold Mariano test for the German-Luxembourg bidding zone. The horizontal model represents model A, while the vertical one represents model B. With a 5% significance level, the null hypothesis is rejected if the p-value is less than 0.05. Based on the legend, shades of green indicate rejection of the null, whereas shades of red indicate failure to reject.
Figure 6: Pairwise DM Test P-Values Across Forecasting Models for German-Luxembourg market
MLP with Reduced Linear (BOA) is statistically more accurate than any other model except BOA all. While MLP with Reduced Linear is the best model according to RMSE, DM results show that this model is not statistically better than MLP with Full Linear, MLP with Reduced Linear (BOA), and BOA all. This may be attributed to the choice of loss function in the DM test. Using a squared loss function might change the results. The red color in LEAR and DNN cells confirms the previous remark about them as the worst-performing models, given that DNN is only statically more accurate than LEAR, and LEAR is not statistically better than any model. Similar results are observed for the Spanish market, where the benchmark models are not statistically more precise than any of the developed models.
Figure 7 illustrates the trade-off between the run time of a model and its accuracy. The best performing model in terms of time is a Full Linear (Reduced Linear for Spain) with a running time of 14 seconds for the whole 730 iterations. This result is consistent with expectations, as linear models typically require the shortest execution time ([a]). MLP with Reduced Linear (BOA) is the slowest model within the Pareto frontier models for both markets. At the same time, it is the model with the best MAE among all models.
While the benchmark models are the worst-performing models in terms of accuracy, their performance in terms of time is much worse. It took LEAR about two hours to train and generate forecasts for the two years of testing, whereas DNN needed about five hours in the German-Luxembourg market and more than three hours for Spain.
Figure 7: Runtine and MAE of considered forecasting models with efficient Pareto Frontier.
6 Conclusion
We propose a novel approach that combines linear and nonlinear mechanisms to develop a forecasting model for electricity prices with high accuracy and computational efficiency. We implemented a main architecture that has two main components: a skip connection from input to output and a nonlinear part connecting the input to the output through a hidden layer. In addition, we developed an alternative architecture that has only the skip connection without the nonlinear part.
We adopt an ensemble method, following the fully adaptive BOA approach, to aggregate individual forecast models and produce more accurate predictions. Moreover, a novel nonparametric partial online learning method is introduced to reduce the computational cost, such that the optimal weights of the previous iteration are employed as initialization of the current iteration.
The models’ performance is evaluated against two state-of-the-art benchmarks: LEAR and DNN models, using German-Luxembourg and Spanish bidding zones data, which are among the largest European markets. The measures of accuracy demonstrate the superiority of all the proposed models over the benchmark models with a 12% reduction in RMSE and 18% decrease in MAE for the German-Luxembourg market and a decline of 13% in RMSE and 15% in MAE for Spain. The results of the Diebold Mariano test are consistent with measures of accuracy metrics. Both LEAR and DNN fail to outperform any of the proposed models.
In addition to their strong predictive accuracy, the suggested models achieve an outstanding runtime. It took the fast models only a few seconds to fit and generate the forecast, while even the slowest and most accurate models complete training and predictions within a few minutes. By contrast, the runtimes of benchmark models are substantially higher, such that each one of them took a couple of hours. This suggests that the proposed online learning approach was very effective in reducing computational time.
In future work, we aim to extend those models to generate probabilistic forecasts due to the limitations of point forecasts in a world full of uncertainties, especially with the large integration of renewables in the grid. Additionally, we plan to explore more advanced neural network architectures such as RNN and LSTM, while carefully monitoring the computational time.
Acknowledgements
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) project numbers 520388526 to BE and FZ.
Declaration on AI assisted tools
During the preparation of this work, the authors used ChatGPT (OpenAI) to improve the readability and clarity of the manuscript and enhance code quality. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
[…]
Appendix
Figure 8: Time Series of Day-Ahead Electricity Prices in Spain
Figure 9: Day-Ahead Forecasts of Load and Renewable Generation, and Commodity Prices in the Spanish Electricity Market
Figure 10: Hourly RMSE of selected models for Spanish bidding zone
Figure 11: Pairwise DM Test P-Values Across Forecasting Models for Spanish market